# CrossGen: Learning and Generating Cross Fields for Quad Meshing

### v1

@@ Abstract.
Cross fields play a critical role in various geometry processing tasks, especially for quad mesh generation. Existing methods for cross field generation often struggle to balance computational efficiency with generation quality, using slow per-shape optimization. We introduce CrossGen, a novel framework that supports both feed-forward prediction and latent generative modeling of cross fields for quad meshing by unifying geometry and cross field representations within a joint latent space. Our method enables extremely fast computation of high-quality cross fields of general input shapes, typically within one second without per-shape optimization. Our method assumes a point-sampled surface, or called a point-cloud surface, as input, so we can accommodate various different surface representations by a straightforward point sampling process. Using an auto-encoder network architecture, we encode input point-cloud surfaces into a sparse voxel grid with fine-grained latent spaces, which are decoded into both SDF-based surface geometry and cross fields (see the teaser figure). We also contribute a dataset of models with both high-quality signed distance fields (SDFs) representations and their corresponding cross fields, and use it to train our network. Once trained, the network is capable of computing a cross field of an input surface in a feed-forward manner, ensuring high geometric fidelity, noise resilience, and rapid inference. Furthermore, leveraging the same unified latent representation, we incorporate a diffusion model for computing cross fields of new shapes generated from partial input, such as sketches. To demonstrate its practical applications, we validate CrossGen on the quad mesh generation task for a large variety of surface shapes. Experimental results demonstrate that CrossGen generalizes well across diverse shapes and consistently yields high-fidelity cross fields, thus facilitating the generation of high-quality quad meshes.
@@
copyright: noneccs: Computing methodologies Shape analysisccs: Computing methodologies Mesh geometry models
Figure 1. Results produced by our method, CrossGen, that predicts cross fields on given shapes for quad meshing in a feed-forward manner without per shape optimization. (a) Input shapes as point cloud surfaces; (b) Cross fields generated by CrossGen; and (c) The resulting quad meshes. CrossGen demonstrates significant advantages in terms of efficiency and generalizability across various shape types, enabling fast high-quality quad mesh generation for downstream applications.
@@ 1. Introduction
Cross fields are fundamental in various geometry processing tasks, such as quadrilateral (quad) mesh generation (Jakob et al., 2015; Huang et al., 2018; Pietroni et al., 2021; Bommes et al., 2009; Dielen et al., 2021; Bommes et al., 2013b; Zhang et al., 2020), surface subdivision (Loop and Schaefer, 2008; Shen et al., 2014), layout generation (Campen et al., 2012), and shape parameterization (T-spline design) (Campen and Zorin, 2017), as they provide a powerful way to capture the underlying structure and flow of surface geometry. By aligning with principal curvature directions and conforming to sharp feature edges or the boundaries of open surfaces, cross fields effectively capture the natural flow of the underlying geometry, leading to more coherent and visually consistent results.
@@ 2.1. Cross Field Generation
Most of the above methods prioritize cross field smoothness at the expense of accurate alignment with principal curvature directions. Ideally, a high-quality cross field should strike a balance between smoothness and faithful alignment to principal curvature directions or sharp features. To this end, NeurCross (Dong et al., 2025) jointly optimizes the cross field and a neural signed distance field (SDF), where the zero-level set of the SDF serves as a proxy for the input surface. NeurCross (Dong et al., 2025) outperforms existing methods in terms of robustness to surface noise and geometric fluctuations, and alignment with curvature directions and sharp feature curves. However, its optimization process is computationally intensive, limiting its applicability in real-time or interactive scenarios. Moreover, all of these methods require per-shape optimization and thus not generalizable to new shapes. Our method, in contrast, is data-driven and therefore generalizable. It can generate cross fields for new shapes without the need for per-shape optimization.
@@ 2.2. Cross Field Applications
[…]
@@
3. Method
Our goal is to efficiently generate a high-quality cross field of a given input point-cloud surface. We propose CrossGen based on an autoencoder architecture with novel designs of both encoder and decoder to enhance the generalization to unseen shapes. Fig. 3 shows the pipeline of CrossGen. In the following we describe the nework architecture, loss functions, and training dataset.
@@ 3.1. Network Architecture
We adopt point clouds as the input surface format, which conveniently accommodates raw point-cloud surface as we as other surface representations, since any other surface representation can be converted to a point-cloud representation with a simple uniform sampling process. The resulting input to the encoder $\mathcal{E}$ is a 3D point set. Let the input point set be denoted by $\mathcal{P}=\{\boldsymbol{p}\mid\boldsymbol{p}\in\mathbb{R}^{3}\}$, where each point $\boldsymbol{p}$ is assumed to be associated with its corresponding surface unit normal vector $\boldsymbol{n}_{\boldsymbol{p}}\in\mathbb{R}^{3}$ as an input feature. These surface normal vectors can be extracted from the original continuous surface if provided as input, or estimated using Xu et al. (2023) for a raw point-cloud surface. These points are initially quantized into the vertices of a $256^{3}$ high-resolution cubic grid to ensure that the geometric details are well captured. This quantized data is subsequently encoded into 16-dimensional latent embeddings by a linear projection layer and fed into the encoder $\mathcal{E}$, which uses a sparse convolutional neural network (CNN) architecture (Choy et al., 2019) – a crucial choice for efficiently processing high-resolution 3D data while significantly reducing computational overhead.
The encoder consists of a sequence of convolutional layers and residual blocks that progressively downsample the input representation. As data propagates through the network, with the successive spatial resolutions of the feature maps being $256^{3}$, $128^{3}$, $64^{3}$, $32^{3}$, and $16^{3}$, and corresponding feature dimensions being 16, 32, 64, 128, and 128. Each downsampling step uses a single convolutional layer to ensure a small receptive field and encourage the encoding of local geometric structures rather than global semantic features. At the network bottleneck, i.e. latent space, the 3D shape is compactly represented as a set of latent patch descriptors =$\mathcal{F}=\{\boldsymbol{f}\mid\boldsymbol{f}\in\mathbb{R}^{m}\}$, where $m$ ($m=128$ by default) denotes the feature dimensionality, as shown in Fig. 3 (b).
@@
The receptive field (RF) of a CNN describes how much context of the input a network can “see” when making predictions at a given location. It defines the geometric context available for encoding the input point cloud surface and predicting the output, i.e. SDFs or cross fields. As shown in the inset for 2D illustration, smaller receptive fields focus on local details, while larger ones capture broader shape context. As discussed in Section 3.1, although shapes from different categories may vary globally, they often share local geometric similarities. To leverage this, we employ a small receptive field in the encoder to focus on local geometry (see Fig. 4). Meanwhile, we adopt a large receptive field in the decoder to aggregate broader context, which promotes smoothness and global coherence in the predicted cross fields.
@@
[…]
@@ 3.2. Loss Function
In our pipeline, both the encoder $\mathcal{E}$ and decoder $\mathcal{D}$ adopt the sparse CNN architecture (Choy et al., 2019) for good computational and memory efficiency. In the encoder $\mathcal{E}$, empty voxels are directly removed during the downsampling process, keeping only the occupied regions for efficient feature extraction. During upsampling by the decoder $\mathcal{D}$, newly generated fine-grained sub-voxels, denoted as $\mathcal{V}$, are retained and marked as occupied only if they intersect with the thin shell region surrounding the zero-level set of the SDF; the rest are discarded.
To supervise voxel occupancy prediction at each resolution level of the output of the decoder $\mathcal{D}$, we introduce an occupancy loss $\mathcal{L}_{\text{o}}$, defined as the discrepancy between the predicted occupancy values $\mathcal{D}_{\text{o}}(\boldsymbol{v})$ and the ground truth occupancy labels $\mathcal{G}_{\text{o}}(\boldsymbol{v})$.
@@
Given a surface point $\boldsymbol{p}$, the cross field branch $\mathcal{M}_{\text{cf}}$ predicts one direction vector of the cross field at $\boldsymbol{p}$, denoted by $\mathcal{M}_{\text{cf}}(\boldsymbol{p})$. This predicted vector $\mathcal{M}_{\text{cf}}(\boldsymbol{p})$ needs to be constrained to lie within the tangent plane at $\boldsymbol{p}$. Hence, we project $\mathcal{M}\text{cf}(\boldsymbol{p})$ onto the tangent plane defined by $\boldsymbol{p}$ and its corresponding GT unit normal vector $\boldsymbol{n}_{\boldsymbol{p}}$.
@@
As illustrated in the inset figure, we consider a local region centered at $\boldsymbol{p}$. Using the unit normal $\boldsymbol{n}_{\boldsymbol{p}}$, we project the predicted vector onto the tangent plane, resulting in $\bar{\boldsymbol{\alpha}}_{\boldsymbol{p}}$, a direction vector that lies within the tangent plane at $\boldsymbol{p}$. That is
@@
[…]
@@
To prevent high variance and stabilize the latent space =$\mathcal{F}=\{\boldsymbol{f}\mid\boldsymbol{f}\in\mathbb{R}^{128}\}$, we incorporate a KL divergence regularization term $D_{\text{KL}}$ (Kingma and Welling, 2013), following (Mo et al., 2019; Rombach et al., 2022). This constrains the latent space distribution $\mathcal{N}(\mathbb{E}(f),\sigma^{2}(f))$ toward a standard normal distribution $\mathcal{N}(0,1)$. Both empirical evidence and established theoretical results (Van Den Oord et al., 2017; Yan et al., 2022; Mittal et al., 2022; Kingma and Welling, 2013) demonstrate that this distributional constraint ensures a topologically coherent and smooth latent space (Mo et al., 2019; Rombach et al., 2022), substantially reducing information entropy during continuous shape parametrization transformations. Hence, we define the latent space regularization as follows:
@@ 3.3. Dataset for Training
To enhance model resilience and promote orientation-invariant cross field predictions, the 1,700 distinct shapes in the dataset are augmented with random rotations, yielding over 10,000 shapes, and all the shapes are normalized to the cube $[-0.5,0.5]^{3}$ to ensure consistent scaling across the shapes. Fig. 6 presents examples of cross field and SDF surfaces included in our dataset. This training dataset is partitioned using a 9:1 train-test split ratio.
@@ 4. Experiments
MIQ  (Bommes et al., 2009) 0.15 265.68 1201.38
Power Fields  (Knöppel et al., 2013) 0.10 3.13 6.31
PolyVectors  (Diamanti et al., 2014) 0.08 1.41 2.17
IM  (Jakob et al., 2015) 0.06 2.13 5.28
QuadriFlow  (Huang et al., 2018) 0.06 2.37 5.58
NeurCross  (Dong et al., 2025) 0.04 360.58 912.46
CrossGen (Ours) 0.04 0.067 0.079
@@
[…]
Figure 8. Quad meshes generated by six baseline methods and our CrossGen. Benefiting from large-scale data-driven learning, our method produces smoother and more regular quad meshes compared to all other approaches.
@@ 4.1. Comparisons
Tab. 1 presents a quantitative comparison between our method and the baseline approaches. On the test data, our method and NeurCross (Dong et al., 2025) achieve the lowest angular error (AE) compared to the principal curvature directions, outperforming all other baselines. Benefiting from the direct inference design, CrossGen achieves the fastest average runtime, outperforming NeurCross (Dong et al., 2025) by approximately 5382 times on the full test data.
Note that only NeurCross (Dong et al., 2025) and our method are GPU-accelerated; all other baselines run on the CPU. To further evaluate performance under high-resolution inputs, we tested 50 randomly selected shapes from the test set, each with over 100,000 faces. As shown in Tab. 1, our method maintains a consistent runtime regardless of input resolution, demonstrating resolution-invariant efficiency. In contrast, all CPU-based methods show significantly increased runtime with higher resolution, while NeurCross (Dong et al., 2025), despite running on the GPU, is over 11,550 times slower than CrossGen and becomes increasingly constrained by computational overhead at higher resolutions. Fig. 7 presents a visual comparison of the cross fields generated by various methods.
@@
Quad Mesh Evaluation
To evaluate the effectiveness of our method in downstream applications, we apply it to quadrilateral meshing tasks and compare the resulting mesh quality with optimization-based approaches in Tab. 2. Except for IM (Jakob et al., 2015) and QuadriFlow (Huang et al., 2018), all methods employ the same quad mesh extraction technique, libQEx (Ebke et al., 2013). Note that our method assumes a point-cloud surface as input, whereas the existing quad mesh extraction methods requires a triangle mesh as the base surface for parameterization and quad extraction. Since all the baseline methods operate on ground-truth triangle meshes, to evaluate our method comprehensively, we consider two settings in Tab. 2: (1) We test our model using the same ground-truth mesh as used the baseline methods as the base surface for querying the cross field and quad meshing, denoted as CrossGen. (2) Starting from point cloud input, we use the SDF branch to reconstruct a surface via Marching Cubes (Lorensen and Cline, 1987), which is then used as the base surface for quad meshing, denoted as CrossGen.
Tab. 2 shows the quantitative results comparing with baseline methods. Our method ranks second only to NeurCross across multiple evaluation metrics, outperforming all other baselines. We note that since NeurCross is used as the expert model to generate the GT data for training our CrossGen network, its performance is expected to be an upper bound for CrossGen’s performance. Besides, while both our method and NeurCross produce a higher number of singularities compared to MIQ (Bommes et al., 2009), our method achieves lower Jacobian Ratio (JR) values, indicating that MIQ (Bommes et al., 2009) tends to generate more distorted quadrilaterals. Fig. 8 presents a visual comparison, where our method produces high-quality and structurally consistent quadrilateral meshes, with results that are overall better or comparable to those of all the baseline methods. In addition, the setting CrossGen, which reconstructs the surface from point cloud inputs via the SDF branch, achieves results comparable to CrossGen, which uses the GT mesh as input. A visual comparison of the two settings is shown in Fig. 9. These results highlight the effectiveness of our SDF branch in producing high-quality surface geometry suitable for downstream quad meshing tasks.

### v2

@@ Abstract.
Cross fields play a critical role in various geometry processing tasks, especially for quad mesh generation. Existing methods for cross field generation often struggle to balance computational efficiency with generation quality, using slow per-shape optimization. We introduce CrossGen, a novel framework that supports both feed-forward prediction and latent generative modeling of cross fields for quad meshing by unifying geometry and cross field representations within a joint latent space. Our method enables extremely fast computation of high-quality cross fields of general input shapes, typically within one second without per-shape optimization. Our method assumes a point-sampled surface, also called a point-cloud surface, as input, so we can accommodate various surface representations by a straightforward point sampling process. Using an auto-encoder network architecture, we encode input point-cloud surfaces into a sparse voxel grid with fine-grained latent spaces, which are decoded into both SDF-based surface geometry and cross fields (see the teaser figure). We also contribute a dataset of models with both high-quality signed distance fields (SDFs) representations and their corresponding cross fields, and use it to train our network. Once trained, the network is capable of computing a cross field of an input surface in a feed-forward manner, ensuring high geometric fidelity, noise resilience, and rapid inference. Furthermore, leveraging the same unified latent representation, we incorporate a diffusion model for computing cross fields of new shapes generated from partial input, such as sketches. To demonstrate its practical applications, we validate CrossGen on the quad mesh generation task for a large variety of surface shapes. Experimental results demonstrate that CrossGen generalizes well across diverse shapes and consistently yields high-fidelity cross fields, thus facilitating the generation of high-quality quad meshes.
@@
journal: TOGcopyright: ccjournal: TOGjournalyear: 2025journalvolume: 44journalnumber: 6publicationmonth: 12doi: 10.1145/3763299ccs: Computing methodologies Shape analysisccs: Computing methodologies Mesh geometry models
Figure 1. Results produced by our method, CrossGen, that predicts cross fields on given shapes for quad meshing in a feed-forward manner without per shape optimization. (a) Input shapes as point cloud surfaces; (b) Cross fields generated by CrossGen; and (c) The resulting quad meshes. CrossGen demonstrates significant advantages in terms of efficiency and generalizability across various shape types, enabling fast, high-quality quad mesh generation for downstream applications.
@@ 1. Introduction
Cross fields are fundamental in various geometry processing tasks, such as quadrilateral (quad) mesh generation (Dielen et al., 2021; Bommes et al., 2013b; Zhang et al., 2020; Viertel and Osting, 2019; Beaufort et al., 2017; Brandt et al., 2018; Palmer et al., 2024), surface subdivision (Loop and Schaefer, 2008; Shen et al., 2014), layout generation (Campen et al., 2012), and shape parameterization (T-spline design) (Campen and Zorin, 2017), as they provide a powerful way to capture the underlying structure and flow of surface geometry. By aligning with principal curvature directions and conforming to sharp feature edges or the boundaries of open surfaces, cross fields effectively capture the natural flow of the underlying geometry, leading to more coherent and visually consistent results.
@@ 2.1. Cross Field Generation
Most existing methods prioritize the smoothness of cross fields but fail to accurately align with curvature directions. Ideally, a high-quality cross field should balance smoothness with faithful alignment to principal curvature directions or sharp geometric features. To this end, NeurCross (Dong et al., 2025) jointly optimizes the cross field and a neural signed distance field (SDF), where the zero-level set of the SDF serves as a proxy for the input surface. NeurCross (Dong et al., 2025) outperforms existing methods in terms of robustness to surface noise and geometric fluctuations, and alignment with curvature directions and sharp feature curves. However, its optimization process is computationally intensive, limiting its applicability in real-time or interactive scenarios. Moreover, these methods rely on per-shape optimization, which is often time-consuming. In contrast, our approach is data-driven and enables direct inference on unseen shapes without requiring additional optimization.
@@ 2.2. Cross Field Applications
3. Method
Our goal is to efficiently generate a high-quality cross field of a given input point-cloud surface. We propose CrossGen based on an autoencoder architecture with novel designs of both encoder and decoder to enhance the generalization to unseen shapes. Fig. 3 shows the pipeline of CrossGen. In the following, we describe the network architecture, loss functions, and training dataset.
@@ 3.1. Network Architecture
We adopt point clouds as the input surface format, which conveniently accommodates raw point-cloud surface as well as other surface representations, since any other surface representation can be converted to a point-cloud representation with a simple uniform sampling process. The resulting input to the encoder $\mathcal{E}$ is a 3D point set normalized to the unit cube. Let the input point set be denoted by $\mathcal{P}=\{\boldsymbol{p}\mid\boldsymbol{p}\in\mathbb{R}^{3}\}$, where each point $\boldsymbol{p}$ is assumed to be associated with its corresponding surface unit normal vector $\boldsymbol{n}_{\boldsymbol{p}}\in\mathbb{R}^{3}$ as an input feature. These surface normal vectors can be extracted from the original continuous surface if provided as input, or estimated using Xu et al. (2023) for a raw point-cloud surface. The point set $\mathcal{P}$ is first quantized onto the vertices of a high-resolution cubic grid $\Omega$, with each point $\boldsymbol{p}$ mapped to its nearest voxel vertex. Voxels containing at least one quantized point are designated as active, whereas all others are regarded as empty and omitted from subsequent sparse convolution operations. To preserve fine geometric details, a sufficiently high grid resolution $\Omega$ is desirable. However, increasing the resolution inevitably incurs substantial memory and computational overhead. In practice, we set $\Omega=256^{3}$, which offers a balanced compromise between fidelity and efficiency. This quantized data is subsequently encoded into 16-dimensional latent embeddings by a linear projection layer and fed into the encoder $\mathcal{E}$, which uses a sparse convolutional neural network (CNN) architecture (Choy et al., 2019) – a crucial choice for efficiently processing high-resolution 3D data while significantly reducing computational overhead.
The encoder consists of a sequence of convolutional layers and residual blocks that progressively downsample the input representation. As data propagates through the network, with the successive spatial resolutions of the feature maps being $256^{3}$, $128^{3}$, $64^{3}$, $32^{3}$, and $16^{3}$, and corresponding feature dimensions being 16, 32, 64, 128, and 128. Each downsampling step uses a single convolutional layer to ensure a small receptive field and encourage the encoding of local geometric structures rather than global semantic features. At the network bottleneck, i.e. latent space, the 3D shape is compactly represented as a set of latent patch descriptors $\mathcal{F}=\{\boldsymbol{f}\mid\boldsymbol{f}\in\mathbb{R}^{m}\}$, where $m$ ($m=128$ by default) denotes the feature dimensionality, as shown in Fig. 3 (b).
@@
The receptive field (RF) of a CNN describes how much context of the input a network can “see” when making predictions at a given location. It defines the geometric context available for encoding the input point cloud surface and predicting the output, i.e. SDFs or cross fields. As shown in the inset for 2D illustration, smaller receptive fields focus on local details, while larger ones capture broader shape context. As discussed in Section 3.1, although shapes from different categories may vary globally, they often share local geometric similarities. To leverage this, we employ a small receptive field in the encoder to focus on local geometry (see Fig. 4). Meanwhile, we adopt a large receptive field in the decoder to aggregate broader context, which promotes smoothness and global coherence in the predicted cross fields. In other words, locality is governed by the receptive field of each voxel in the sparse CNN, rather than by explicit patching or segmentation. The encoder’s use of small kernels and shallow layers ensures that voxel features retain fine-grained geometric detail. Moreover, because sparse convolutions operate over a global voxel grid, receptive fields of adjacent voxels naturally overlap, promoting smooth regional transitions. This convolutional hierarchy implicitly aggregates neighboring features, eliminating the need for explicit fusion operations such as averaging.
@@
Figure 5. Overview of our SDF and cross field modules. Both modules are implemented as MLPs that decode the high-resolution feature grid into SDFs and cross field, respectively. $\mathcal{*}$ represents trilinear interpolation operation.
@@ 3.2. Loss Function
Since raw point clouds are often noisy, sparse, and incomplete, direct occupancy estimation can be unreliable; therefore, we incorporate an occupancy loss into our framework. This loss also acts as a regularizer during training, encouraging the network to learn stronger geometric priors across shapes and thereby maintain local geometric consistency.
In our implementation, both the encoder $\mathcal{E}$ and decoder $\mathcal{D}$ adopt the sparse CNN architecture (Choy et al., 2019) for good computational and memory efficiency. In the encoder $\mathcal{E}$, empty voxels are directly removed during the downsampling process, keeping only the occupied regions for efficient feature extraction. During upsampling by the decoder $\mathcal{D}$, newly generated fine-grained sub-voxels, denoted as $\mathcal{V}$, are retained and marked as occupied only if they intersect with the thin shell region surrounding the zero-level set of the SDF; the rest are discarded. To supervise voxel occupancy prediction at each resolution level of the output of the decoder $\mathcal{D}$, we introduce an occupancy loss $\mathcal{L}_{\text{o}}$, defined as the discrepancy between the predicted occupancy values $\mathcal{D}_{\text{o}}(\boldsymbol{v})$ and the ground truth occupancy labels $\mathcal{G}_{\text{o}}(\boldsymbol{v})$.
@@
Given a surface point $\boldsymbol{p}$, the cross field branch $\mathcal{M}_{\text{cf}}$ predicts one direction vector of the cross field at $\boldsymbol{p}$, denoted by $\mathcal{M}_{\text{cf}}(\boldsymbol{p})$. This predicted vector $\mathcal{M}_{\text{cf}}(\boldsymbol{p})$ needs to be constrained to lie within the tangent plane at $\boldsymbol{p}$. Hence, we project $\mathcal{M}_{\text{cf}}(\boldsymbol{p})$ onto the tangent plane defined by $\boldsymbol{p}$ and its corresponding GT unit normal vector $\boldsymbol{n}_{\boldsymbol{p}}$. As illustrated in the inset figure, we consider a local region centered at $\boldsymbol{p}$. Using the unit normal $\boldsymbol{n}_{\boldsymbol{p}}$, we project the predicted vector onto the tangent plane, resulting in $\bar{\boldsymbol{\alpha}}_{\boldsymbol{p}}$, a direction vector that lies within the tangent plane at $\boldsymbol{p}$. That is
@@
where $\left\|\cdot\right\|$ denotes the absolute value.
@@
To prevent high variance and stabilize the latent space $\mathcal{F}=\{\boldsymbol{f}\mid\boldsymbol{f}\in\mathbb{R}^{128}\}$, we incorporate a KL divergence regularization term $D_{\text{KL}}$ (Kingma and Welling, 2013), following (Mo et al., 2019; Rombach et al., 2022). This constrains the latent space distribution $\mathcal{N}(\mathbb{E}(f),\sigma^{2}(f))$ toward a standard normal distribution $\mathcal{N}(0,1)$. Both empirical evidence and established theoretical results (Van Den Oord et al., 2017; Yan et al., 2022; Mittal et al., 2022; Kingma and Welling, 2013) demonstrate that this distributional constraint ensures a topologically coherent and smooth latent space (Mo et al., 2019; Rombach et al., 2022), substantially reducing information entropy during continuous shape parametrization transformations. Hence, we define the latent space regularization as follows:
@@ 3.3. Dataset for Training
To enhance model resilience and promote orientation-invariant cross field predictions, the 1,700 distinct shapes in the dataset are augmented with random rotations, where each shape is rotated by a random angle about a random axis, yielding over 10,000 shapes. All the shapes are normalized to the cube $0.5,0.5]^{3}$ to ensure consistent scaling across the shapes. Fig. 6 presents examples of cross field and SDF surfaces included in our dataset. The dataset is partitioned at the distinct shape level using a 9:1 train–test split.
@@ 4. Experiments
MIQ  (Bommes et al., 2009) 0.17 265.68 1201.38
Power Fields  (Knöppel et al., 2013) 0.11 3.13 6.31
PolyVectors  (Diamanti et al., 2014) 0.10 1.41 2.17
IM  (Jakob et al., 2015) 0.07 2.13 5.28
QuadriFlow  (Huang et al., 2018) 0.07 2.37 5.58
NeurCross  (Dong et al., 2025) 0.05 360.58 912.46
CrossGen (Ours) 0.05 0.067 0.079
@@ 4.1. Comparisons
Evaluating cross fields is inherently challenging, as principal curvature directions may become ambiguous in noisy or flat regions. To ensure rigorous and reliable assessment, we therefore select 50 smooth, noise-free shapes with well-defined principal directions, which minimizes ambiguity and provides a trustworthy reference for evaluation. It should be noted that these 50 smooth, noise-free models are used exclusively for computing the angular error reported in Tab. 1, which presents a quantitative comparison between our method and the baseline approaches. On the test data, our method and NeurCross (Dong et al., 2025) achieve the lowest angular error (AE) compared to the principal curvature directions, outperforming all other baselines. Benefiting from the direct inference design, CrossGen achieves the fastest average runtime, outperforming NeurCross (Dong et al., 2025) by approximately 5382 times on the full test data.
Note that only NeurCross (Dong et al., 2025) and our method are GPU-accelerated; all other baselines run on the CPU. To further evaluate performance under high-resolution inputs, we tested 50 randomly selected shapes from the test set, each with over 100,000 faces. These 50 models are different from the 50 smooth, noise-free models used to compute the angular error. As shown in Tab. 1, our method maintains a consistent runtime regardless of input resolution. In contrast, all CPU-based methods show significantly increased runtime with higher resolution, while NeurCross (Dong et al., 2025), despite running on the GPU, is over 11,550 times slower than CrossGen and becomes increasingly constrained by computational overhead at higher resolutions. Fig. 7 presents a visual comparison of the cross fields generated by various methods.
@@
Figure 8. Quad meshes generated by six baseline methods and our CrossGen. The red points mark the locations of singularities, where “# of Sings” denotes the number of singular points in each quad mesh. Benefiting from large-scale data-driven learning, our method produces smoother and more regular quad meshes compared to all other approaches.
[…]
@@
Quad Mesh Evaluation
To evaluate the effectiveness of our method in downstream applications, we apply it to quadrilateral meshing tasks and compare the resulting mesh quality with optimization-based approaches in Tab. 2. Except for IM (Jakob et al., 2015) and QuadriFlow (Huang et al., 2018), all methods employ the same quad mesh extraction technique, libQEx (Ebke et al., 2013). Note that our method assumes a point-cloud surface as input, whereas the existing quad mesh extraction methods requires a triangle mesh as the base surface for parameterization and quad extraction. Since all the baseline methods operate on ground-truth triangle meshes, to evaluate our method comprehensively, we consider two settings in Tab. 2: (1) We test our model using the same ground-truth mesh as the base surface for querying the cross field and quad meshing, denoted as CrossGen. (2) Starting from point cloud input, we use the SDF branch to reconstruct a surface via Marching Cubes (Lorensen and Cline, 1987), which is then used as the base surface for quad meshing, denoted as CrossGen.
Tab. 2 shows the quantitative results comparing with baseline methods. Our method ranks second only to NeurCross across multiple evaluation metrics, outperforming all other baselines. We note that since NeurCross is used as the expert model to generate the GT data for training our CrossGen network, its performance is expected to be an upper bound for CrossGen’s performance. Furthermore, although both our method and NeurCross (Dong et al., 2025) produce a greater number of singularities than MIQ (Bommes et al., 2009), they achieve higher Jacobian Ratio (JR) values, indicating that MIQ (Bommes et al., 2009) tends to generate more distorted quadrilaterals. Fig. 8 presents a visual comparison, where our method produces high-quality and structurally consistent quadrilateral meshes, with results that are overall better or comparable to those of all the baseline methods. In Fig. 8, we also highlight the locations of singularities. The distribution of singular points in the quad mesh generated by our CrossGen is similar to that produced by NeurCross (Dong et al., 2025).
In addition, the setting of CrossGen, which reconstructs the surface from point cloud inputs via the SDF branch, achieves results comparable to CrossGen, which uses the GT mesh as input. A visual comparison of the two settings is shown in Fig. 9. These results highlight the effectiveness of our SDF branch in producing high-quality surface geometry suitable for downstream quad meshing tasks.
[…]
Comparison with Point2Quad
@@
[…]
Figure 12. Quad meshes generated by our method under varying input conditions, where the normal of our input point cloud is obtained using Xu et al. (2023) method. (a) the original point cloud with 150K points; (b) a point cloud with missing regions (highlighted in light violet); (c) an uneven point cloud input; and (d) Point cloud inputs with varying levels of sparsity.